OpenStudy (anonymous):

The legs of a right triangle measure 5 inches and 7 inches. If theta is the angle between the 5-inch leg and the hypotenuse, sin theta = _____. A. .07 B. .58 C. .71 D. .81

OpenStudy (anonymous):



OpenStudy (anonymous):

hypotenuse is \(\sqrt{5^2+7^2}=\sqrt{25+49}=\sqrt{74}\) and the sine is \(\frac{7}{\sqrt{74}}\)

OpenStudy (anonymous):

7/√74=.813 so D :) thanks!!

OpenStudy (anonymous):

yw